OpenStudy (anonymous):

whatat is the distribution property of 33+45

OpenStudy (anonymous):

http://math.about.com/od/algebra/a/distributive.htm

OpenStudy (anonymous):

@blairebradshaw

OpenStudy (judygreeneyes):

To figure that out you need to find the factors they have in common. 33 and 45 can be factored as: 33 = 3 * 11 45 = 3 * 3 * 5 They have a 3 in common, so you can write 33 + 45 as 33 + 45 = 3(11 + 15)

OpenStudy (judygreeneyes):

:)

OpenStudy (anonymous):

@judygreeneyes may I have a medal?